OpenStudy (anonymous):

An apple sells for 25 cents and a peach sells for 15 cents. a total of 10 pieces of fruit were sold for a total cost of $2.10(210 cents). How many apples were sold?

OpenStudy (anonymous):

let's consider that x apples were sold and y peches were sold... Since total of 10 fruits were sold, no. of apples + no. of peaches = 10. let x apples were sold. Guess how many peaches were sold then?? that is 10 - x. price of x apples = x * 25 price of ( 10 - x ) peaches = (10 - x) * 15 . now, 25x + (10 - x) . 15 = 210. do the calculation. the answer comes out to be 6.